Circular No. 7767

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE B. Barris, University of Hawaii, on behalf of the High-Z Supernova

Search Team (cf. IAUC 6160, 6646, 7312, 7516, 7745, 7755), reports the discovery of several apparent supernovae on I-band images taken with the Canada-France-Hawaii Telescope (+ 12K Mosaic in queue mode); the supernovae were identified by subtracting images taken on Sept. 11 and Oct. 9 and 13 from those taken Nov. 12, and confirmed on images obtained by P. Challis and E. Martin (Subaru telescope + SUPRIME) on Nov. 18, 19, and 23. Spectral observations were made with the Keck II telescope (+ ESI) by R. Ellis and M. Sullivan (Nov. 13) and by A. V. Filippenko and R. Chornock (Nov. 18 and 19). SNe 2001hn and 2001ho were discovered in the Sept./Oct. data (cf. IAUC 7745) and are now extremely faint. In the table below, I_{D} is the discovery I magnitude, and I_{L} is the magnitude determined on 2001 Nov. 18 (or Oct. 9 for SNe 2001hn and 2001ho). The tabulated redshifts are for the host galaxy’s emission. Each apparent supernova is within 2′′ of its host galaxy’s center.

SN  2001 UT  α_{2000}  δ_{2000}  I_{D}  I_{L}  z
2001hi Nov. 12 2^{h}26^{m}50.^{s}04 + 0^{o}33′03′′.4 24.3 24.1
2001hj Nov. 12 2 27 30.80 + 0 24 17.0 24.7 24.6
2001hk Nov. 12 2 27 50.29 + 0 36 59.8 23.5 23.5
2001hl Nov. 12 2 28 26.89 + 0 36 19.4 24.4 24.5
2001hm Nov. 12 2 28 32.02 + 0 30 03.1 24.4 24.4
2001hn Sept. 24 2 29 00.26 + 0 43 20.6 23.3 24.3
2001ho Sept. 24 2 29 10.66 + 0 21 36.9 22.9 23.3 0.401
2001hp Nov. 12 2 29 17.18 + 0 25 48.0 23.3 23.6 0.335
2001hq Nov. 12 4 37 36.24 − 1 24 42.9 24.6 23.8
2001hr Nov. 12 4 38 20.55 − 1 33 57.4 23.9 24.2 0.515
2001hs Nov. 12 4 39 22.39 − 1 32 51.4 24.2 23.6 0.80
2001ht Nov. 12 4 39 57.62 − 1 37 41.7 24.4 23.9
2001hu Nov. 12 7 50 35.90 + 9 58 14.2 24.4 24.1 0.88
2001hv Nov. 12 8 47 40.10 +44 23 59.7 24.0 23.9 0.47
2001hw Nov. 12 8 48 15.08 +44 03 38.5 23.7 23.8
2001hx Nov. 12 8 49 24.61 +44 02 22.4 23.6 23.4 0.80
2001hy Nov. 12 8 49 45.85 +44 15 31.8 23.7 23.6 0.81
2001hz Nov. 12 8 49 50.35 +44 22 21.8 23.6 23.6
2001ia Nov. 12 8 49 52.54 +44 11 37.8 24.0 24.3
2001 December 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green